Solve by using the quadratic formula x^2=-7x-3

Answer by checkley71(8403)

X^2=-7X-3
X^2+7X+3=0
X=(-B+-SQRT[B^2-4AC])/2A
X=(-7+-SQRT[49-4*1*3])/2*1
X=(-7+-SQRT[49-12])/2
X=(-7+-SQRT37)/2
X=(-7+-6.08)/2
X=(-7+6.08)/2
X=-.92/2
X=-.46 SOLUTION
X=(-7-6.08)/2
X=-13.08/2
X=-6.54 SOLUTION